OpenStudy (anonymous):

Solve.... 5w + 6 = 9 – w What is w?

OpenStudy (anonymous):

5w+6=9-w 5w+w=9-6 6w=3 w=1/2

jimthompson5910 (jim_thompson5910):

5w + 6 = 9 – w 5w + 6-6 = 9 – w-6 5w = 3 - w 5w +w= 3 - w +w 6w = 3 6w/6 = 3/6 w = 1/2

OpenStudy (anonymous):

thankyou :)

jimthompson5910 (jim_thompson5910):

yw